### v2

# Simplicity Prevails: The Emergence of Generalizable AIGI Detection in Visual Foundation Models

Abstract.
Specialized detectors for AI-generated images (AIGI) often achieve near-perfect accuracy on curated benchmarks, yet their performance degrades substantially in realistic, in-the-wild scenarios. In this work, we show that frozen features from modern Vision Foundation Models (VFMs), combined with a lightweight classifier, form a remarkably strong baseline for generalizable AIGI detection. Using representative modern encoders, including Perception Encoder, MetaCLIP 2, and DINOv3, we conduct a comprehensive evaluation across standard benchmarks, recent unseen generators, and challenging in-the-wild distributions. Across these settings, this simple baseline consistently matches or outperforms recent specialized detectors, with particularly large gains in realistic scenarios.
We further investigate why this simple setup is so effective. Our analyses provide converging evidence that the strong forensic separability of modern VFMs is closely related to their exposure to synthetic web content during pre-training. In Vision-Language Models, this manifests as semantic alignment with forgery-related concepts, while in Self-Supervised Learning models it appears as implicit discrimination of generative distributions. Although a fully controlled pre-training study is beyond the scope of this work, multiple complementary analyses support this interpretation. We also identify important limitations. While modern VFMs are highly effective for global AIGI detection, they remain vulnerable to severe transmission degradation and perform poorly on pure VAE reconstruction and localized editing. Overall, our results suggest that progress in generalizable AIGI detection may depend more on preserving and leveraging strong pretrained representations than on increasingly complex task-specific forensic designs.
AI-generated image detection, multimedia forensics, vision foundation models
copyright: noneconference: Proceedings of the 34th ACM International Conference on Multimedia; October 2026; Amsterdam, The Netherlandsbooktitle: Proceedings of the 34th ACM International Conference on Multimedia (ACM $M$ ’26), October 2026, Amsterdam, The Netherlandsjournalyear: 2026doi: TBDisbn: TBDccs: Computing methodologies Computer visionccs: Computing methodologies Artificial intelligenceccs: Information systems Multimedia content analysis
1. Introduction
“One thing that should be learned from the bitter lesson is the great power of general purpose methods, of methods that continue to scale with increased computation even as the available computation becomes very great.”
— Rich Sutton, The Bitter Lesson (Sutton, 2019)
The rapid evolution of generative models, such as Midjourney, Stable Diffusion (Rombach et al., 2021) and Nano Banana(Team et al., 2023), has ushered in a new era of content creation, synthesizing photorealistic images that challenge the boundaries of visual authenticity. While empowering creativity, this technological leap simultaneously introduces profound threats to information integrity, fueling the proliferation of misinformation. In response, the forensics community has largely favored a specialized approach: crafting detectors with increasingly complex module tailored to specific artifacts, such as frequency anomalies or noise residuals (Wang et al., 2020; Ju et al., 2022). While these specialized detectors achieve near-perfect accuracy on curated benchmarks, they often suffer from a dramatic performance collapse in realistic scenarios. Recent studies, such as the Chameleon benchmark (Yan et al., 2024a), reveal that detectors excelling in controlled environments frequently degrade to 60%–70% accuracy when deployed ‘in-the-wild’. This fragility suggests that relying on hand-crafted inductive biases may be a dead end in the face of rapidly evolving generative distributions.
Figure 1. The Surge of Generative Data in Web Corpora. We track the number of indexed URLs from major open-source AI generation platforms (Civitai and Liblib) within Common Crawl snapshots from 2022 to 2025.
Echoing Sutton’s “Bitter Lesson”, we revisit this trend from a different perspective: rather than asking how to design increasingly specialized forensic modules, we ask how far one can go by directly leveraging the frozen representations of modern Vision Foundation Models (VFMs). We show that a simple linear classifier, trained on top of frozen features from recent encoders such as Perception Encoder (PE) (Bolya et al., 2025), MetaCLIP 2 (Chuang et al., 2025), and DINOv3 (Siméoni et al., 2025), provides a remarkably strong baseline for generalizable AIGI detection. We define modern VFMs as the latest generation of encoders trained on large-scale and evolving web corpora. Our evaluation spans three distinct protocols: standard benchmarks (e.g., GenImage (Zhu et al., 2023)), datasets from the latest unseen generators (e.g., AIGIHolmes (Zhou et al., 2025b), AIGI-Now (Chen et al., 2025a)), and challenging in-the-wild distributions (e.g., Chameleon (Yan et al., 2024a), WildRF (Cavia et al., 2024)). Across all settings, this simple baseline consistently matches or outperforms recent specialized detectors, with the largest gains appearing in the most challenging in-the-wild scenarios.
We further investigate why such a simple setup is so effective. Rather than attributing the gains to forensic-specific architectural innovation, we argue that they are closely related to an emergent property of large-scale pre-training on evolving web data. As visualized in Figure 1, our analysis of the Common Crawl index reveals a rapid increase in generative content from major communities such as Civitai and Liblib starting in 2023. This trend suggests that modern VFMs are increasingly likely to encounter synthetic content during pre-training, and may therefore internalize useful cues for distinguishing generated images from real ones. We characterize this capability through two distinct manifestations of data exposure. For Vision-Language Models, the co-occurrence of synthetic images and textual descriptions can lead to explicit concept injection, where synthetic visuals become aligned with high-level forgery-related concepts. For Self-Supervised Learning (SSL) models such as DINOv3, the capability appears instead as implicit distribution fitting, where the model captures low-level regularities associated with the generative manifold through pre-training data exposure.
At the same time, our analysis also clarifies the boundaries of this paradigm. We find that modern VFMs remain blind to pure reconstruction artifacts and struggle with localized editing, while also suffering from noticeable degradation under aggressive real-world transmission and screen recapture. These findings suggest that large-scale pre-training has substantially improved the generalizability of global detection for fully synthetic content, but that the localization of fine-grained manipulation remains an open challenge. Complementing these observations, our experiments on backbone replacement and lightweight LoRA fine-tuning suggest that the dominant factor behind performance is still the strength of the pretrained representation itself, while adaptation mainly serves to better exploit this foundation. Therefore, rather than viewing increasingly specialized detector design as the default path forward, we argue that future progress in AI forensics may depend more on how to preserve, harness, and refine the evolving representations of foundation models for fine-grained forensic reasoning.
In summary, our main contributions are:
  • We show that frozen features from modern Vision Foundation Models, combined with a lightweight classifier, constitute a remarkably strong baseline for generalizable AIGI detection. Across standard benchmarks, recent unseen generators, and challenging in-the-wild distributions, this simple setup consistently matches or outperforms recent specialized detectors, with especially large gains in realistic scenarios.
  • We provide converging evidence that this capability is closely related to pre-training data exposure rather than forensic-specific architectural design. In particular, we identify two complementary manifestations: semantic alignment with forgery-related concepts in Vision-Language Models, and implicit discrimination of generative distributions in Self-Supervised Learning models.
  • We validate the “Bitter Lesson” in AI forensics and delineate the boundaries of this paradigm. Through rigorous ablation, we demonstrate that attaching complex forensic heads or applying fine-tuning (e.g., LoRA) actively degrades the generic representations of VFMs. While these frozen generic features solve global detection, they remain blind to pure VAE reconstruction and localized editing, urging future research to harness rather than over-engineer foundation models.
Table 1. Performance on GenImage Benchmark. All detectors are trained on Stable Diffusion v1.4 and evaluated on unseen generators. Accuracy is averaged over real and fake classes. Best results in bold.
Method ADM BigGAN Midjourney VQDM GLIDE SD-v1.4 SD-v1.5 Wukong Avg.
Modern VFM Baselines (Ours)
MetaCLIP-Linear 0.549 0.524 0.839 0.653 0.653 0.984 0.981 0.942 0.766
MetaCLIP2-Linear 0.690 0.816 0.959 0.819 0.887 0.993 0.991 0.980 0.892
SigLIP-Linear 0.740 0.658 0.812 0.857 0.916 0.955 0.954 0.918 0.851
SigLIP2-Linear 0.870 0.924 0.879 0.954 0.951 0.995 0.995 0.994 0.945
PE-CLIP-Linear 0.712 0.963 0.901 0.959 0.972 0.999 0.999 0.999 0.938
DINOv2-Linear 0.651 0.889 0.811 0.757 0.852 0.967 0.961 0.929 0.852
DINOv3-Linear 0.849 0.991 0.934 0.992 0.963 0.998 0.996 0.994 0.964
Competitor Methods
CNNSpot (Wang et al., 2020) 0.507 0.530 0.610 0.501 0.522 0.998 0.996 0.985 0.706
FreqNet (Tan et al., 2024a) 0.608 0.939 0.849 0.626 0.858 0.950 0.949 0.937 0.840
Gram-Net (Liu et al., 2020) 0.576 0.591 0.665 0.528 0.747 0.989 0.988 0.960 0.756
NPR (Tan et al., 2024b) 0.619 0.849 0.850 0.567 0.844 0.996 0.994 0.984 0.838
UnivFD (Ojha et al., 2023) 0.571 0.839 0.795 0.644 0.842 0.959 0.958 0.926 0.817
SAFE (Li et al., 2025b) 0.600 0.880 0.829 0.708 0.836 0.964 0.967 0.903 0.836
LaDeDa (Cavia et al., 2024) 0.512 0.583 0.595 0.513 0.585 0.876 0.871 0.820 0.670
Effort (Yan et al., 2024b) 0.704 0.749 0.756 0.824 0.807 0.874 0.873 0.867 0.807
DDA (Chen et al., 2025b) 0.880 0.741 0.960 0.719 0.862 0.986 0.985 0.986 0.890
OMAT (Zhou et al., 2025a) 0.837 0.973 0.903 0.954 0.974 0.974 0.973 0.975 0.946
AIDE (Yan et al., 2024a) 0.602 0.648 0.813 0.694 0.514 0.959 0.957 0.954 0.768
2. Related Works
The development of AI-generated image (AIGI) detection has undergone a significant paradigm shift, evolving from hand-crafted artifact analysis to the adaptation of large-scale foundation models.
Early Artifact-Based Detection. Initial forensic methods focused on identifying the low-level imperfections inherent to early generative architectures. Researchers found that upsampling operations in GANs and CNNs often leave distinct footprints, such as checkerboard artifacts in the pixel space (Odena et al., 2016) or spectral anomalies in the frequency domain (Frank et al., 2020). Others exploited inconsistencies in color statistics (McCloskey and Albright, 2019) or noise residuals (Cozzolino and Verdoliva, 2019) to distinguish synthetic content. While effective against specific generators, these hand-crafted features proved brittle against the rapid evolution of generative models, particularly with the advent of Diffusion Models which exhibit fundamentally different artifact patterns.
Data-Driven Specialized Detectors. With the dominance of Diffusion Models, the focus shifted from detecting CNN-specific upsampling artifacts (e.g., checkerboard patterns in GANs (Wang et al., 2020)) to identifying the unique traces of the diffusion process. Researchers proposed reconstructing input images to isolate generative errors: DIRE (Wang et al., 2023) leverages the reconstruction residual from a pre-trained diffusion model as a forensic signal, while DRCT (Chen et al., 2024) refines this by analyzing the discrepancies between real-real and fake-fake reconstruction pairs. Others focused on improving generalization: SAFE (Li et al., 2025b) introduces artifact-preserving augmentations to decouple semantic content from forensic traces. notably, DDA (Chen et al., 2025b) targets the shared VAE decoder inherent to Latent Diffusion Models, explicitly aligning the detector with VAE reconstruction patterns to achieve broader generalization across different LDM-based generators.
The Foundation Model Era. The introduction of UnivFD  (Ojha et al., 2023) marked a pivotal turning point. Ojha et al. revealed that training a linear layer on top of the frozen feature space of a pre-trained Vision-Language Model (specifically CLIP (Radford et al., 2021)) yields significantly better generalization than training CNNs from scratch. This discovery spurred a wave of research leveraging Vision Foundation Models (VFMs) as backbones. Subsequent works, such as Effort(Yan et al., 2024b), AIDE (Yan et al., 2024a), OMAT(Zhou et al., 2025a) and DDA(Chen et al., 2025b), have explored various strategies to adapt CLIP or DINOv2 (Oquab et al., 2023) for forensics, including prompt tuning, adapter modules, fusing frequency-domain information and training hard samples. These methods currently represent the state-of-the-art, yet as our experiments show, they still struggle to maintain robustness in unconstrained, in-the-wild scenarios compared to the raw capabilities of the latest VFMs. (See Appendix for a comprehensive survey of the legacy backbones still prevalent in these recent methods.)
3. Simplicity Prevails: Benchmarking Modern VFMs
To empirically validate our hypothesis that the generalization capability of AIGI detection stems from the scale of pre-training data rather than complex architectural designs, we conduct a comprehensive comparative analysis. We pit simple linear classifiers trained on modern Vision Foundation Models (VFMs) against a wide array of state-of-the-art specialized detectors. Our evaluation protocol is designed to be rigorous and progressively challenging, spanning three distinct scenarios: standard academic benchmarks, realistic in-the-wild distributions, and unseen next-generation generative models.
3.1. Experimental Setup
Evaluation Benchmarks. To rigorously assess generalization, we organize our evaluation into three progressively challenging categories. (1) Standard Benchmarks: We use GenImage (Zhu et al., 2023), a widely adopted benchmark comprising images from 8 generators (e.g., Stable Diffusion, Midjourney). Following standard protocols (Ojha et al., 2023), we use the Stable Diffusion v1.4 subset for training and the remaining subsets for testing. (2) In-the-Wild Datasets: We evaluate on Chameleon (Yan et al., 2024a), WildRF (Cavia et al., 2024), SocialRF and CommunityAI (Li et al., 2025c). These datasets are collected from social media and internet forums, featuring diverse, unconstrained post-processing and unknown generative sources, representing a realistic detection scenario. (3) Unseen Generators: We employ AIGIHolmes (Zhou et al., 2025b) and AIGI-Now (Chen et al., 2025a), recent benchmarks containing images from state-of-the-art generators released after 2024, including closed-source generators like Nano Banana, GPT4o and FLUX-Pro. These serve as a strict test for generalization to unseen distributions.
Specialized Detectors. We compare against a comprehensive suite of state-of-the-art forensic methods, spanning three categories: (1) Specialized Detectors: Artifact-based methods like CNNSpot (Wang et al., 2020), FreqNet (Tan et al., 2024a), and NPR (Tan et al., 2024b), as well as recent VFM-based adapters such as UnivFD (Ojha et al., 2023), OMAT (Zhou et al., 2025a), Effort (Yan et al., 2024b) and Dual-Data-Alignment (DDA)(Chen et al., 2025b). (2) Former VFM Baselines: To explicitly evaluate the impact of pre-training data evolution, we include earlier generations of foundation models, such as the original OpenAI CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023), Meta CLIP (Xu et al., 2023) and DINOv2 (Oquab et al., 2023). For DDA, we utilize the official pre-trained weights provided by the authors, as its core contribution involves a specialized training pipeline with VAE-reconstructed data alignment. All other methods are trained on GenImage SDv1.4 training set for fair comparison.
Modern VFM Baselines. To test our “Simplicity Prevails” hypothesis, we select a representative set of modern Vision Foundation Models as frozen feature extractors. These include Vision-Language Models (SigLIP2 (Tschannen et al., 2025), MetaCLIP 2 (Chuang et al., 2025), Perception Encoder (Bolya et al., 2025)) and Self-Supervised Models (DINOv3 (Siméoni et al., 2025)). We attach a simple linear layer to the pooled output features of these backbones. Detailed specifications of model architectures and pre-training datasets are provided in Appendix.
Table 2. Performance on In-the-Wild Benchmarks. Evaluation on Chameleon, WildRF, SocialRF, and CommunityAI datasets. Accuracy is averaged over real and fake classes. Best results in bold.
Method Chameleon WildRF SocialRF CommunityAI Avg.
Real Fake Avg. Real Fake Avg. Real Fake Avg. Real Fake Avg.
Modern VFM Baselines (Ours)
MetaCLIP-Linear 0.373 0.914 0.644 0.461 0.923 0.692 0.409 0.866 0.638 0.353 0.933 0.643 0.654
MetaCLIP2-Linear 0.948 0.913 0.930 0.478 0.979 0.728 0.659 0.940 0.800 0.926 0.954 0.940 0.842
SigLIP-Linear 0.480 0.732 0.606 0.383 0.897 0.640 0.549 0.613 0.581 0.370 0.857 0.614 0.610
SigLIP2-Linear 0.884 0.833 0.859 0.597 0.984 0.790 0.744 0.866 0.805 0.826 0.905 0.866 0.822
PE-CLIP-Linear 0.970 0.948 0.959 0.679 0.994 0.836 0.751 0.970 0.861 0.966 0.975 0.971 0.899
DINOv2-Linear 0.628 0.580 0.608 0.643 0.772 0.705 0.603 0.695 0.649 0.606 0.562 0.583 0.636
DINOv3-Linear 0.933 0.895 0.914 0.948 0.975 0.961 0.937 0.948 0.943 0.949 0.946 0.948 0.940
Competitor Methods
CNNSpot 0.979 0.128 0.554 0.959 0.290 0.625 0.588 0.541 0.565 0.969 0.112 0.541 0.571
FreqNet 0.985 0.090 0.538 0.731 0.559 0.645 0.544 0.553 0.549 0.977 0.128 0.553 0.571
Gram-Net 0.992 0.044 0.518 0.947 0.205 0.576 0.531 0.523 0.527 0.985 0.061 0.523 0.536
NPR 0.999 0.046 0.523 0.980 0.243 0.612 0.593 0.537 0.565 0.998 0.076 0.537 0.560
UnivFD 0.763 0.441 0.602 0.693 0.637 0.665 0.563 0.637 0.600 0.778 0.495 0.636 0.617
SAFE 0.993 0.046 0.520 0.918 0.309 0.613 0.564 0.532 0.548 0.984 0.080 0.532 0.556
LaDeDa 0.994 0.015 0.504 0.988 0.119 0.554 0.542 0.506 0.524 0.986 0.026 0.506 0.523
Effort 0.394 0.782 0.588 0.179 0.955 0.567 0.513 0.533 0.523 0.225 0.840 0.533 0.553
DDA 0.940 0.708 0.824 0.899 0.908 0.904 0.818 0.846 0.832 0.968 0.725 0.847 0.850
OMAT 0.899 0.359 0.629 0.633 0.715 0.674 0.581 0.633 0.607 0.853 0.414 0.634 0.636
AIDE 0.944 0.203 0.574 0.973 0.195 0.584 0.578 0.541 0.560 0.990 0.093 0.542 0.565
Implementation Details. All our VFM baselines are trained solely on the GenImage (SD v1.4) training set. We keep the backbone completely frozen and only update the linear head. We use the AdamW optimizer with a learning rate of $1e^{-3}$ and a batch size of 128 for 2 epoch. Images are resized and center-cropped to the native resolution of each model without any additional data augmentation. Notably, PE uses a ViT-L/14 backbone at 336px resolution, which is comparable in scale to several strong VFM-based baselines and recent detectors built on CLIP-L style encoders.
3.2. Performance on Standard Benchmarks
We first establish the efficacy of our simple baselines on the standard GenImage benchmark in Table 1. Despite the simplicity of the linear probe, modern VFMs achieve state-of-the-art performance. DINOv3-Linear reaches the highest average accuracy of 96.5%, surpassing the best specialized detector OMAT (94.6%) and significantly outperforming legacy baselines. Notably, we observe a substantial performance leap across VFM generations: DINOv3 improves upon DINOv2 by over 11%, and MetaCLIP-2 boosts accuracy by 12.6% compared to its predecessor MetaCLIP. This trend highlights that forensic discriminability is not static but scales with the quality and data volume of the foundation model. Furthermore, while specialized detectors often overfit to the training source, modern VFMs demonstrate robust generalization across diverse generative architectures, confirming that their representations are inherently forensic-ready without the need for complex auxiliary modules.
3.3. The Collapse of SOTA in the Wild
While standard benchmarks provide a controlled environment, real-world deployment involves diverse, unconstrained data distributions. To evaluate this, we test on four challenging in-the-wild datasets: Chameleon, WildRF, Social-RF, and CommunityAI. The results, summarized in Table 2, reveal a stark contrast. Most specialized detectors suffer a catastrophic performance collapse. Methods like NPR, LaDeDa, and even recent techniques like OMAT degrade to near-random performance, primarily due to a failure to recognize diverse fake samples. The only exception among specialized methods is DDA, which maintains a respectable average accuracy of 85.0%, due to its alignment with the robust VAE decoder patterns shared across latent diffusion models by training on DINOv2 with VAE reconstruct data.
However, modern VFM baselines decisively outperform all competitors. DINOv3-Linear achieves an average accuracy of 94.0%, surpassing DDA by nearly 10% and traditional detectors by over 30%. Crucially, we observe a massive performance gap between modern and legacy VFMs: DINOv3 outperforms its predecessor DINOv2 by a staggering 30.4%, and MetaCLIP-2 surpasses MetaCLIP by 18.8%. This confirms that earlier foundation models lacked the necessary data exposure to handle in-the-wild shifts, whereas modern iterations have internalized these distributions during pre-training.
Table 3. Generalization on AIGIHolmes. Evaluation on advanced auto-regressive and diffusion-transformer generative models. Accuracy is averaged over real and fake classes. Best results in bold.
Detector FLUX Infinity Janus Janus-Pro-1B Janus-Pro-7B LlamaGen PixArt-XL SD3.5-L Show-o VAR Avg
Modern VFM Baselines (Ours)
MetaCLIP-Linear 0.951 0.978 0.736 0.950 0.950 0.970 0.970 0.867 0.972 0.614 0.896
MetaCLIP2-Linear 0.987 0.990 0.839 0.959 0.928 0.989 0.986 0.956 0.985 0.802 0.942
SigLIP-Linear 0.888 0.918 0.889 0.928 0.926 0.913 0.896 0.826 0.868 0.834 0.889
SigLIP2-Linear 0.957 0.994 0.990 0.993 0.989 0.991 0.994 0.914 0.992 0.913 0.973
PE-CLIP-Linear 0.968 0.999 0.945 0.995 0.996 1.000 1.000 0.943 0.999 0.935 0.978
DINOv3-Linear 0.933 0.998 0.996 0.995 0.986 0.999 0.999 0.891 0.997 0.922 0.972
Competitor Methods
CNNSpot 0.626 0.589 0.508 0.531 0.522 0.604 0.668 0.568 0.608 0.504 0.573
FreqNet 0.894 0.924 0.443 0.439 0.433 0.875 0.885 0.894 0.910 0.688 0.738
Gram-Net 0.696 0.604 0.491 0.491 0.491 0.627 0.627 0.701 0.771 0.557 0.606
NPR 0.968 0.983 0.495 0.495 0.495 0.947 0.899 0.937 0.988 0.730 0.794
LaDeDa 0.682 0.664 0.498 0.497 0.497 0.810 0.641 0.695 0.755 0.724 0.646
UnivFD 0.868 0.898 0.575 0.674 0.575 0.915 0.925 0.898 0.911 0.597 0.784
SAFE 0.932 0.965 0.482 0.482 0.483 0.913 0.896 0.940 0.962 0.911 0.796
Effort-AIGI 0.794 0.804 0.483 0.650 0.569 0.798 0.804 0.765 0.793 0.764 0.722
DDA 0.972 0.989 0.985 0.993 0.987 0.993 0.994 0.970 0.948 0.804 0.963
OMAT 0.947 0.955 0.651 0.756 0.641 0.967 0.969 0.957 0.969 0.959 0.877
AIDE 0.944 0.987 0.912 0.989 0.978 0.994 0.986 0.994 0.980 0.936 0.970
Table 4. Generalization on AIGI-Now. Evaluation on 9 open-source and closed-source generative models. Accuracy is averaged over real and fake classes. Best results in bold.
Detector FLUX-dev FLUX-kera FLUX-kontext FLUX-pro gpt4o jimeng keling minimax Nano Avg.
pix sem pix sem pix sem pix sem pix sem pix sem pix sem pix sem pix sem
Modern VFM Baselines (Ours)
MetaCLIP-Linear 0.948 0.965 0.956 0.913 0.708 0.808 0.894 0.943 0.898 0.882 0.846 0.829 0.963 0.939 0.889 0.877 0.917 0.884 0.892
MetaCLIP2-Linear 0.979 0.941 0.963 0.896 0.799 0.811 0.976 0.892 0.943 0.888 0.965 0.825 0.970 0.902 0.942 0.850 0.965 0.819 0.907
SigLIP-Linear 0.838 0.898 0.840 0.891 0.700 0.811 0.836 0.911 0.827 0.897 0.827 0.890 0.863 0.937 0.848 0.863 0.796 0.867 0.852
SigLIP2-Linear 0.947 0.882 0.883 0.697 0.776 0.678 0.888 0.885 0.936 0.790 0.831 0.845 0.941 0.867 0.850 0.688 0.895 0.882 0.843
PE-CLIP-Linear 0.977 0.959 0.918 0.762 0.830 0.774 0.873 0.943 0.863 0.924 0.915 0.921 0.939 0.916 0.865 0.748 0.971 0.936 0.891
DINOv3-Linear 0.944 0.962 0.846 0.811 0.730 0.756 0.813 0.948 0.898 0.960 0.824 0.940 0.884 0.913 0.727 0.784 0.898 0.922 0.864
Competitor Methods
CNNSpot 0.919 0.500 0.550 0.500 0.843 0.502 0.535 0.500 0.990 0.501 0.523 0.500 0.973 0.501 0.603 0.504 0.985 0.499 0.635
FreqNet 0.875 0.468 0.697 0.443 0.769 0.479 0.492 0.501 0.923 0.510 0.459 0.530 0.922 0.543 0.824 0.511 0.907 0.487 0.622
Gram-Net 0.933 0.528 0.667 0.522 0.864 0.555 0.608 0.569 0.763 0.508 0.508 0.503 0.955 0.554 0.717 0.521 0.905 0.528 0.651
NPR 0.944 0.500 0.508 0.500 0.785 0.502 0.502 0.502 0.966 0.500 0.497 0.501 0.957 0.500 0.548 0.500 0.930 0.500 0.619
LaDeDa 0.586 0.498 0.497 0.497 0.560 0.502 0.496 0.495 0.745 0.499 0.495 0.497 0.661 0.501 0.509 0.502 0.766 0.505 0.546
UnivFD 0.542 0.579 0.514 0.538 0.492 0.545 0.516 0.544 0.529 0.532 0.475 0.504 0.631 0.539 0.501 0.539 0.501 0.516 0.531
SAFE 0.903 0.490 0.532 0.492 0.831 0.494 0.521 0.486 0.977 0.488 0.509 0.486 0.960 0.487 0.590 0.491 0.961 0.484 0.621
Effort-AIGI 0.789 0.679 0.796 0.669 0.728 0.610 0.688 0.690 0.753 0.580 0.522 0.555 0.782 0.677 0.772 0.687 0.796 0.657 0.690
DDA 0.916 0.512 0.594 0.499 0.827 0.529 0.766 0.550 0.923 0.654 0.870 0.654 0.961 0.646 0.833 0.505 0.816 0.562 0.695
OMAT 0.911 0.475 0.649 0.469 0.847 0.507 0.591 0.515 0.744 0.452 0.491 0.465 0.936 0.526 0.699 0.467 0.891 0.468 0.615
AIDE 0.991 0.590 0.504 0.569 0.979 0.806 0.601 0.538 0.747 0.518 0.639 0.514 0.982 0.554 0.514 0.541 0.989 0.518 0.672
3.4. Generalization to State-of-the-Art Generators
A critical question remains: do these models truly learn generalized forensic concepts, or do they merely memorize pre-training patterns? We investigate this via AIGIHolmes and AIGI-Now, which challenge detectors across two fronts:
  • AIGIHolmes: Features recent Auto-Regressive (AR) models (e.g., LlamaGen, VAR) and Diffusion Transformers (e.g., FLUX), whose mechanisms differ fundamentally from older UNet-based models (e.g., SD-v1.4).
  • AIGI-Now: Contains closed-source APIs (e.g., GPT-4o, FLUX-Pro) unseen during VFM pre-training. And disentangles evaluation into two subsets: Pixel-artifact (pix), which isolates low-level generative traces by strictly aligning image formats; and Semantic (sem), which applies aggressive degradations to obliterate low-level artifacts, forcing detectors to rely solely on high-level semantic anomalies.
Results. As shown in Tables 3 and 4, modern VFMs demonstrate exceptional transferability. On AIGIHolmes, PE-CLIP and DINOv3 achieve average accuracies of 97.8% and 97.2%, maintaining robust detection even on fundamentally distinct AR models (e.g., VAR).
On AIGI-Now, MetaCLIP2 leads with 90.7%. Crucially, VFMs excel across both the pix and heavily degraded sem subsets, proving they capture a dual-level (structural and semantic) notion of artificiality. Conversely, specialized detectors like CNNSpot collapse to near-random guessing ($\sim$50%) on the sem splits. This confirms that modern VFMs learn generalized, robust forensic concepts that extend far beyond specific generators or brittle low-level artifacts.
Table 5. Comparison of Text–Image Similarities on In-the-Wild Dataset
Method Chameleon SocialRF Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) modern design 0.628 portrait 0.345 forged 0.332 urban 0.253 Siglip (2023.3.27) unaltered 0.318 original 0.212 edited 0.232 original 0.222 Siglip2 (2025.2.21) genuine 0.385 urban 0.250 portrait 0.202 vintage 0.191 Meta CLIP (2023.9.28) AI generated 0.678 deepfake 0.091 AI generated 0.902 original 0.024 Meta CLIP-2 (2025.7.29) AI generated 0.828 deepfake 0.064 AI generated 0.924 deepfake 0.038 PE (2025.4.17) AI generated 0.861 deepfake 0.021 AI generated 0.943 deepfake 0.031 Method CommunityAI Midjourney-CC Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) portrait 0.346 nature 0.326 urban 0.332 midjourney_images 0.260 Siglip (2023.3.27) AIGIBench 0.283 real 0.281 fake 0.278 original 0.246 Siglip2 (2025.2.21) urban 0.209 portrait 0.201 urban 0.212 portrait 0.194 Meta CLIP (2023.9.28) AI generated 0.726 deepfake 0.086 midjourney_images 0.604 AI generated 0.284 Meta CLIP-2 (2025.7.29) AI generated 0.858 deepfake 0.041 midjourney_images 0.621 AI generated 0.308 PE (2025.4.17) AI generated 0.878 CommunityAI 0.029 midjourney_images 0.722 AI generated 0.218
4. Analysis: The Mechanisms of Emergence
The strong performance of simple probes on modern VFMs raises a central question: does this capability primarily come from forensic-specific architectural choices, or from properties already present in large-scale pretrained representations? Motivated by the rapid growth of synthetic content on the web (Figure 1), we investigate whether pre-training data exposure is an important factor behind this phenomenon. Because fully controlled retraining of proprietary billion-parameter models is computationally infeasible, we do not attempt to make a definitive causal claim. Instead, we use a set of complementary indirect analyses to characterize how this capability may emerge. Taken together, these analyses suggest two main mechanisms: semantic conceptualization in Vision-Language Models and implicit distribution discrimination in Self-Supervised Learning models.
4.1. Mechanism I: Semantic Conceptualization in VLMs
For Vision-Language Models (VLMs), we hypothesize that their capability stems from the contrastive pre-training objective. During training, massive volumes of synthetic images co-occur with metadata or captions containing explicit indicators of their source (e.g., midjourney,AI generated). Consequently, the model internalizes a powerful semantic shortcut: it learns to align the visual features of synthetic content directly with forgery-related textual concepts. To validate this, we conduct a text-image alignment analysis without training any classifier. We probe whether the frozen embedding space of VLMs naturally clusters synthetic images closer to forgery-related prompts. We constructed a comprehensive text pool categorized into three conceptual groups to probe the model’s internal associations:
  • Forgery-Related Concepts: Terms explicitly denoting authenticity or fabrication (e.g., ‘fake’, ‘real’, ‘AI generated’, ‘authentic’, ‘manipulated’, ‘synthetic’).
  • Content-Related Concepts: Neutral descriptions of visual content (e.g., ‘sunset’, ‘landscape’, ‘portrait’, ‘abstract art’, ‘technology’, ‘nature’).
  • Source-Related Concepts: Specific names of generative models or platforms (e.g., ‘GenImage’, ‘ADM’, ‘BigGAN’, ‘glide’, ‘Midjourney’).
We evaluate the cosine similarity on in-the-wild benchmarks and our newly collected Midjourney-CC dataset (3,000 images from reddit.com/r/midjourney, late 2025) to strictly control for data leakage.
Table 5 presents the results of our semantic probing, revealing a striking dichotomy between legacy and modern VLMs. Legacy models like CLIP (2021) and SigLIP (2023) exhibit “forensic blindness”, mapping synthetic images to content terms (e.g., portrait). Notably, even the recently released SigLIP 2 (2025) fails to detect forgery concepts (Top-1: genuine/urban), likely because it relies on the older WebLI dataset (Chen et al., 2022) curated in 2022, prior to the generative explosion. In sharp contrast, modern VLMs trained on recent web crawls (MetaCLIP 2, PE) consistently align fake images with “AI generated”. Crucially, on Midjourney-CC, these models specifically retrieve “midjourney_images”, providing definitive evidence that their capability stems from exposure to recent, platform-specific metadata which older datasets lack.
4.2. Mechanism II: Data-Driven Feature Discrimination in SSL
While VLMs rely on semantic tags, Self-Supervised Learning (SSL) models like DINOv3 lack textual supervision, yet they often outperform VLMs in our benchmarks. We hypothesize that this capability is acquired implicitly through distribution fitting: by training on a massive web corpus mixed with generative content, the model learns to encode the distinct low-level signatures of the generative manifold into its feature space as separable clusters, independent of semantic labels.
To validate that this capability stems from data exposure rather than architectural advantages, we conduct a counterfactual experiment. We employ the identical DINOv3 ViT-7B architecture but vary the pre-training data source: DINOv3-Web (LVD-1689M): Pre-trained on a large-scale web corpus containing 1.6 billion diverse internet images, which naturally includes a significant volume of AIGI. DINOv3-Sat (Sat-493M): Pre-trained on 493 million satellite images, a domain strictly devoid of generative content.
Table 6. Counterfactual Analysis. Comparison of DINOv3 trained on Web Data vs. Satellite Data.
Pre-training Data GenImage Chameleon
Avg. Real Fake Avg.
DINOv3-Web 0.965 0.933 0.895 0.914
DINOv3-Sat 0.706 0.948 0.121 0.535
Table 6 delivers a decisive finding. While the web-trained baseline excels, DINOv3-Sat completely fails on fake images, despite performing well on real ones. This collapse proves that the model classifies unseen fakes as “real” simply because it is not included in pretrained data. The conclusion is the forensic capability of SSL models is not inherent to the architecture or training strategy, but is entirely contingent on exposure to generative data during pre-training.
5. Robustness and Limitations
5.1. Protocol I: Resilience to Common Perturbations
To assess real-world reliability, we evaluate detector robustness against JPEG compression (Quality $\in\{95,\dots,65\}$) and Gaussian Blur ($\sigma\in\{0.5,\dots,2.0\}$). We benchmark modern VFM linear probes against legacy models and specialized detectors across both GenImage and the in-the-wild Chameleon datasets.
Figure 2. Robustness to Common Perturbations. Accuracy trajectories under JPEG compression and Gaussian Blur.
As visualized in Figure 2, the trajectories reveal a stark stratification between legacy VFMs (dashed lines) and modern VFMs (solid lines). Older models not only establish lower baselines but also exhibit severe volatility under perturbations, particularly on Chameleon. This decisively proves that robustness is not an inherent benefit of the linear probing architecture. Instead, modern VFMs exhibit superior resilience because their intrinsic forensic capabilities are derived from massive, inadvertent exposure to diverse synthetic content during web-scale pre-training. By internalizing the generative manifold directly from the messy, unconstrained internet, these models learn robust, high-level features rather than brittle, lab-generated artifacts. While PE still relies somewhat on high-frequency traces susceptible to low-pass filtering (dropping to 77.8% at $\sigma=2.0$), DINOv3 and MetaCLIP2 capture fundamental structural anomalies that inherently resist smoothing. This profound resilience to blur and compression allows our image-trained linear probes (e.g., DINOv3) to achieve generalized SOTA performance on video benchmarks like VidProM and GenVideo via simple frame-level aggregation, decisively outperforming bespoke AI video detectors.
5.2. Protocol II: Real-World Transmission and Recapture
To evaluate deployment reliability, we further assess performance under severe image degradation scenarios, including optical recapture from screens and heavy compression introduced by social media transmission protocols. To evaluate this, we utilize the RRDataset (Li et al., 2025a), measuring performance across three settings: Original (Digital baseline), Redigital (Screen or print recapture), and Transfer (Social media transmission).
Table 7. Robustness Evaluation on RRDataset. We report the accuracy on Real and AI classes separately.
Detector Original (Base) Redigital (Recapture) Transfer (Social App)
Real AI Real AI Real AI
CNNSpot 0.977 0.189 0.986 0.068 0.993 0.065
NPR 0.957 0.749 0.955 0.083 0.998 0.002
SAFE 0.874 0.792 0.928 0.066 0.997 0.009
DDA 0.942 0.895 0.961 0.298 0.933 0.532
SigLIP2-Linear 0.8714 0.931 0.885 0.559 0.490 0.704
MetaCLIP2-Linear 0.784 0.939 0.791 0.719 0.930 0.713
PE-CLIP-Linear 0.925 0.949 0.913 0.548 0.989 0.685
DINOv3-Linear 0.951 0.930 0.964 0.647 0.980 0.712
As detailed in Table 7, specialized detectors suffer a total collapse under transmission and recapture. Methods like SAFE and NPR degrade to near-zero sensitivity on fake images ($1\%$). Even the robust DDA drops to 29.8% on recaptured data. In sharp contrast, modern VFMs maintain robust detection capabilities. MetaCLIP2-Linear leads with $\sim$72% accuracy across both scenarios. Consistent with the blur experiments (Sec. 5.1), PE-CLIP suffers a notable drop (to 54.8%) on recaptured data, confirming its reliance on fine-grained details prone to erasure by low-pass filtering. Conversely, DINOv3 and MetaCLIP2 exhibit superior resilience, suggesting that their learned structural anomalies and semantic concepts persist even through the analog-to-digital bottleneck.
5.3. Protocol III: Robustness to Reconstruction and Editing
While modern VFMs excel at detecting fully generated images, forensic reliability requires handling subtler manipulations: (1) VAE Reconstruction, where a real image is encoded and decoded by a diffusion model’s VAE without semantic modification (simulating deepfake pre-processing); and (2) Local Editing, where only specific regions are inpainted. We evaluate on the DDA-COCO (Chen et al., 2025b) (VAE-reconstructed real images) and BR-Gen  (Cai et al., 2025)(Diffusion-based local editing) datasets.
Table 8. Limitations under Reconstruction and Editing. Detection accuracy on DDA-COCO (VAE-based reconstruction) and BR-Gen (Diffusion-based local editing).
Method DDA-COCO BR-Gen
SDXL SD2 SD3.5L Brush Power SDXL
CNNSpot 0.014 0.017 0.004 0.121 0.091 0.073
DDA 0.949 0.997 0.682 0.648 0.589 0.465
OMAT 0.211 0.304 0.135 0.697 0.715 0.681
Effort 0.511 0.549 0.682 0.801 0.793 0.767
SigLIP2-Linear 0.071 0.079 0.017 0.597 0.623 0.476
MetaCLIP2-Linear 0.057 0.074 0.037 0.544 0.575 0.500
PE-CLIP-Linear 0.066 0.170 0.024 0.564 0.581 0.528
DINOv3-Linear 0.030 0.079 0.004 0.592 0.613 0.450
Table 8 (Left) exposes a critical limitation: modern VFMs are essentially blind to pure VAE reconstruction artifacts. Detection rates plummet to negligible levels, indicating that these models do not perceive the low-level noise footprint of the VAE decoder as an anomaly. Conversely, DDA, which explicitly aligns with VAE reconstruction patterns, maintains robust performance. As shown in Table 8 (Right), VFMs struggle to generalize to localized manipulations (BR-Gen), with performance hovering around 50%–60%. We attribute this to the global pooling mechanism inherent to our linear probing approach: the dominant feature signal from the unaltered “real” regions likely suppresses the subtle forensic traces within the edited mask. In contrast, methods like Effort, designed to amplify anomaly feature, achieve higher accuracy.
5.4. Protocol IV: The Bottleneck of Specialized Architectures
A natural question arises from our primary finding: if modern Vision Foundation Models (VFMs) possess such powerful representations, could their performance be further amplified by attaching State-Of-The-Art specialized forensic architectures? To investigate this, we upgraded recent expert models—Effort (Yan et al., 2024b), AIDE (Yan et al., 2024a), and DDA (Chen et al., 2025b)—by swapping their original legacy backbones with our top-performing modern VFMs (MetaCLIP2, PE, and DINOv3).
Table 9. Impact of Upgrading Specialized Architectures. While modern backbones improve the performance of specialized methods, they still severely underperform the simple frozen linear probe. Original baselines are in italics, best results in bold.
Method & Backbone GenImage Chameleon AIGIHolmes
Effort (Original: CLIP-L) 0.807 0.588 0.722
Effort + MetaCLIP2 0.748 0.720 0.799
Effort + PE 0.856 0.761 0.924
AIDE (Original: ResNet/ConvNeXt) 0.768 0.574 0.970
AIDE + MetaCLIP2 0.834 0.709 0.922
AIDE + PE 0.883 0.914 0.947
DDA (Original: DINOv2) 0.890 0.824 0.963
DDA + MetaCLIP2 0.557 0.563 0.596
DDA + PE 0.534 0.559 0.659
DDA + DINOv3 0.572 0.751 0.713
MetaCLIP2-Linear 0.892 0.930 0.942
PE-CLIP-Linear 0.938 0.959 0.978
DINOv3-Linear 0.964 0.914 0.972
The results in Table 9 expose the Bottleneck of Inductive Bias. While upgrading expert models (e.g., AIDE, Effort) with modern VFMs improves their baseline performance, they still strictly underperform our minimalist linear probe (e.g., AIDE with PE achieves 91.4% on Chameleon, falling short of PE-Linear’s 95.9%). Worse still, rigidly specialized architectures like DDA suffer catastrophic degradation ($\sim$50-60%) when forced onto new generic feature spaces. This reveals that complex inductive biases—such as explicit frequency filtering or strict VAE alignment—actually act as information bottlenecks, inadvertently constraining the raw, universal discriminative power naturally emergent in modern representations.
5.5. Protocol V: The Pitfall of Parameter-Efficient Fine-Tuning
Another prevalent paradigm for adapting foundation models is Parameter-Efficient Fine-Tuning (PEFT). If a simple linear layer suffices, would unfreezing the backbone via Low-Rank Adaptation (LoRA) yield even better task-specific performance? To test this, we applied LoRA with rank $r\in\{4,8\}$ to modern VFMs, fine-tuning them on the GenImage (SD v1.4) training set.
Table 10. LoRA Fine-Tuning vs. Frozen Linear Probe. Unfreezing the backbone via LoRA significantly degrades generalization in the wild. Best results in bold.
Backbone Strategy GenImage Chameleon AIGIHolmes
MetaCLIP2 Linear Probe (Frozen) 0.892 0.930 0.942
LoRA (r=4) 0.734 0.817 0.823
LoRA (r=8) 0.780 0.880 0.896
PE Linear Probe (Frozen) 0.938 0.959 0.978
LoRA (r=4) 0.810 0.719 0.879
LoRA (r=8) 0.761 0.635 0.891
DINOv3 Linear Probe (Frozen) 0.964 0.914 0.972
LoRA (r=4) 0.954 0.803 0.977
LoRA (r=8) 0.928 0.718 0.945
Table 10 highlights the Risk of Modifying Internal Knowledge. Contrary to the intuition that fine-tuning improves task adaptation, applying LoRA actively dismantles generalizability. For instance, LoRA fine-tuning on PE (r=8) causes its in-the-wild detection accuracy (Chameleon) to plummet from 95.9% to a mere 63.5%. We attribute this to manifold distortion and catastrophic forgetting. By unfreezing the backbone to optimize for a single generator (SD v1.4), the model rapidly overfits to a narrow, transient distribution of local artifacts, irrevocably overwriting the broad ”world knowledge” of synthetic anomalies it implicitly derived from massive pre-training data.
6. Conclusion
In this work, we revisit AIGI detection from the perspective of pretrained visual representations. We show that frozen features from modern Vision Foundation Models, combined with a lightweight classifier, form a remarkably strong baseline for generalizable AIGI detection. Across standard benchmarks, in-the-wild datasets, and recent unseen generators, this simple setup consistently matches or outperforms recent specialized detectors. Our analyses further suggest that this capability is closely related to exposure to synthetic web content during pre-training, rather than primarily to forensic-specific architectural design. In VLMs, this appears as semantic alignment with forgery-related concepts, while in SSL models it appears as implicit discrimination of generative distributions. Although fully controlled pre-training ablations are beyond the scope of this work, our evidence consistently supports this interpretation.
At the same time, modern VFMs remain weak on pure VAE reconstruction, localized editing, and severe transmission or recapture. We therefore view frozen modern VFM representations not as a complete solution to multimedia forensics, but as a strong foundation for robust global AIGI detection. More broadly, our findings suggest that future progress may depend less on increasingly specialized detector design, and more on effectively leveraging the evolving representations learned by foundation models.